# Canonical differential equations for the elliptic two-loop five-point integral family relevant to $t\bar{t}+$jet production at leading colour

### v1

@@ IV Conclusion and Outlook
We derived canonical DEs for the two-loop, five-point integral family involving elliptic functions that contributes to the leading colour NNLO QCD corrections to top-pair production with a jet at hadron colliders. Together with the boundary values of ref. [32], this makes it straightforward to write the solution analytically in terms of iterated integrals with closed-form kernels [72]. This result finalises the construction of canonical DEs for all integral families needed for this prediction [31, 32, 33] and opens several interesting research avenues.
@@
The generalisation of the even/odd parity to incorporate nested square roots and the constraints this imposes on Feynman integrals through their DEs are other important insights raised in this work. This aspect is closely tied to the possibility of expressing DEs containing nested square roots in ${\mathrm{d}}\log$ form. Our results lay strong foundation to extend to nested square roots the active study of symbol letters [58, 59, 76, 53, 77, 60, 61, 54, 62, 78, 79, 80, 81, 82, 64, 65, 83, 63, 84, 66, 85, 67, 68, 86], which underpins many bootstrap programs (see ref. [87] for a review).
Another primary challenge is to obtain an analytic representation of the MIs that not only reveals simplifications in the amplitude but also enables efficient numerical evaluation. Recent progress in expressing MIs in terms of independent special functions [88, 89, 90, 91, 92] together with a specialised approach to their evaluation [93, 89] have led to a surge of NNLO QCD predictions for $2\to 3$ LHC processes with massless virtual particles [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109]. Our results set the necessary ground to study the extension of these techniques to elliptic integrals in a context where this is essential to obtain theoretical predictions. In fact, the complexity of the Monte Carlo phase-space integration sets substantially higher efficiency demands for $2\to 3$ processes as compared to lower multiplicities. The extension of the techniques above would thus have a transformative impact, and remove one of the main bottlenecks to improve the predictions for a wide class of processes. More general strategies based on series expansions are also under study for iterated integrals with higher-genus geometries [110, 27, 28, 73, 29], and their potential merits further exploration.
Beyond being a critical step towards achieving NNLO QCD corrections for an important LHC process, our results prove that the methodology under development by the community has reached the maturity to tackle the computation of elliptic Feynman integrals with five-particle kinematics. This milestone is important also in view of other massive $2\to 3$ amplitudes, which are object of active research [111, 45, 112, 113] and play a crucial role in LHC phenomenology. Furthermore, the mathematical challenges addressed in this work are relevant in other researches areas, including gravitational waves [114, 115, 23, 116], string theory [117] and pure mathematics, underscoring the broad impact of our findings.
@@ Acknowledgements.
We thank Simon Badger, Christoph Nega, Vasily Sotnikov and Fabian Wagner for useful discussions, the organisers of the workshop ‘MathemAmplitudes 2023: QFT at the Computational Frontier,’ where this project started, and CERN for hospitality. We are also grateful to Simon Badger, Johannes Henn and Vasily Sotnikov for useful comments on this letter. This project received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101105486 and ERC Starting Grant No. 101040760 FFHiggsTop, and from the Swiss National Science Foundation (SNSF) under the Ambizione grant No. 215960.
Appendix A Identifying the $G_{i}$ with complete elliptic integrals of the third kind
In this appendix we explain how we obtained expressions for the functions $G_{i}$ introduced in section II in terms of complete elliptic integrals of the third kind. The procedure is similar to the one described in ref. [19] around eq. (3.53), with the difference that we cannot always cut all propagators since we are also interested in functions that appear in the inhomogeneous parts of the DEs.
@@ Appendix A Identifying the $G_{i}$ with complete elliptic integrals of the third kind
where $z$ is the Baikov variable associated to the eighth propagator and $y^{2}=(z-e_{1})(z-e_{2})(z-e_{3})(z-e_{4})$ is the elliptic curve. The roots $e_{i}$ are functions of $\vec{x}$ and can be found in ref. [32]. We therefore see why the (properly normalised) periods $\psi_{1}$ and $\psi_{2}$ defined in eq. 7 are solutions to the second-order differential operator $L_{m_{t}^{2}}^{(2)}$ annihilating the first integral at $\epsilon=0$.
Similarly, the transformation to canonical form introduces $G_{1}$ as the solution to a third-order differential operator (related to $L_{36}$ discussed in ref. [32]). In more detail, the DEs in eq. (10) can be rewritten as a third-order differential operator annihilating the third integral, $I_{1,1,0,1,1,1,0,1}^{0,0,0}[(k_{1}+p_{1})^{2}]$, where the term in the square brackets is a numerator under the integral sign. The general solution for this DE is given by
@@
$+c_{2}\left(4m_{t}^{2}\psi_{2}^{2}+G_{1,2}\right)+c_{3}+\ldots\,,$
@@
that evaluates to a complete elliptic integral of the third kind $\Pi$ with argument either $\kappa^{2}$ or $1-\kappa^{2}$ (defined in section I), depending on the integration cycle. For example, we have that
@@
This immediately relates $G_{1}$ to either $\mathrm{\Pi}(\frac{e_{3}-e_{2}}{e_{3}-e_{1}},\kappa^{2})$ or $\mathrm{\Pi}(\frac{e_{1}-e_{2}}{e_{1}-e_{3}},1-\kappa^{2})$ depending on which period is chosen in the semi-simple part $W^{\rm ss}$ discussed in section II.
@@
$+c_{2}\left(8\,G_{2,1}+4\,m_{t}^{2}\,\frac{\psi_{1}}{\Delta_{32}}\right)$
$+c_{3}\left(8\,G_{2,2}+4\,m_{t}^{2}\,\frac{\psi_{2}}{\Delta_{32}}\right)+\ldots\,,$
@@
where $\mathrm{Cut}_{\text{ell.}}$ denotes the cut of the elliptic sector, $f_{1}(z_{7},z_{9})$ is linear in $z_{7}$ and $z_{9}$, and $f_{2}(z_{7},z_{9})$ is quadratic in $z_{7}$ and $z_{9}$. After a change of variables, $\{z_{7},z_{9}\}\to\{x,y\}$, that rationalises both square roots in the denominator, the integral in $x$ can be performed to show that the integrand of eq. 24 has the form
@@
$+\mathrm{E}(k)\frac{(a-1)k^{\prime}-(k-1)a^{\prime}}{2\sqrt{a-1}\sqrt{a}\sqrt{a-k}(k-1)}\,,$
where the prime indicates a partial derivative w.r.t. some kinematic variable. We emphasise that the r.h.s. of eq. 27 does not contain elliptic integrals of the third kind. For example, this relation can be used to determine the prefactor of $\Pi(a,k)$ when the arguments $a$ and $k$ have been identified. Moreover, once the result for one integration cycle is known, the replacement

### v2

@@ IV Conclusion and Outlook
We derived canonical DEs for the two-loop, five-point integral family involving elliptic functions that contributes to the leading color NNLO QCD corrections to top-pair production with a jet at hadron colliders. Together with the boundary values of Ref. [33], this makes it straightforward to write the solution analytically in terms of iterated integrals with closed-form kernels [35]. This result finalizes the construction of canonical DEs for all integral families needed for this prediction [32, 33, 34] and opens several interesting research avenues.
@@
The generalization of the even/odd parity to incorporate nested square roots and the constraints this imposes on Feynman integrals through their DEs are other important insights raised in this work. This aspect is closely tied to the possibility of expressing DEs containing nested square roots in ${\mathrm{d}}\log$ form. Our results lay a strong foundation to extend to nested square roots the active study of symbol letters [61, 62, 78, 56, 79, 63, 64, 57, 65, 80, 81, 82, 83, 84, 67, 68, 85, 66, 86, 69, 87, 70, 71, 88], which underpins many bootstrap programs (see Ref. [89] for a review).
Another primary challenge is to obtain an analytic representation of the MIs that not only reveals simplifications in the amplitude but also enables efficient numerical evaluation. As of now, the best available approach to evaluate the MIs considered here is that of generalized series expansions [90] applied to non-canonical DEs [33, 34]; while flexible, this method suffers from slow evaluation times w.r.t. what is typically needed for Monte-Carlo integration. Recent progress in expressing MIs in terms of independent special functions [91, 92, 93, 94, 95] together with a specialized approach to their evaluation [96, 92] have led to a surge of NNLO QCD predictions for $2\to 3$ LHC processes with massless virtual particles [97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113]. Our results set the necessary ground to study the extension of these techniques to elliptic integrals in a context where this is essential to obtain theoretical predictions. In fact, the complexity of the Monte Carlo phase-space integration sets substantially higher efficiency demands for $2\to 3$ processes as compared to lower multiplicities. The extension of the techniques above would thus have a transformative impact, and remove one of the main bottlenecks to improve the predictions for a wide class of processes. More general strategies based on series expansions are also under study for iterated integrals with higher-genus geometries [48, 27, 28, 75, 30], and their potential merits further exploration.
Beyond being a critical step towards achieving NNLO QCD corrections for an important LHC process, our results prove that the methodology under development by the community has reached the maturity to tackle the computation of elliptic Feynman integrals with five-particle kinematics. This milestone is important also in view of other massive $2\to 3$ amplitudes, which are the object of active research [114, 47, 115, 116, 117] and play a crucial role in LHC phenomenology. Furthermore, the mathematical challenges addressed in this work are relevant in other research areas, including gravitational waves [118, 119, 23, 120], string theory [121] and pure mathematics, underscoring the broad impact of our findings.
@@ Acknowledgements.
We thank S. Badger, C. Nega, V. Sotnikov, L. Tancredi and F. Wagner for useful discussions, the organizers of ‘MathemAmplitudes 2023,’ and CERN for hospitality. We are grateful to S. Badger, J. Henn and V. Sotnikov for useful comments on the draft. This project received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101105486 and ERC Starting Grant No. 101040760 FFHiggsTop, and from the Swiss National Science Foundation under the Ambizione grant No. 215960.
Appendix A Explicit expressions for the $G_{i}$’s
In this appendix we explain how we obtained expressions for the functions $G_{i}$ appearing in the canonical DEs in terms of complete elliptic integrals of the third kind. The procedure is similar to the one described in Ref. [19] around Eq. (3.53), with the difference that we cannot always cut all propagators since we are also interested in functions that appear in the inhomogeneous parts of the DEs.
@@ Appendix A Explicit expressions for the $G_{i}$’s
where $z$ is the Baikov variable associated to the eighth propagator and $y^{2}=(z-e_{1})(z-e_{2})(z-e_{3})(z-e_{4})$ is the elliptic curve. The roots $e_{i}$ are functions of $\vec{x}$ and can be found in Ref. [33]. We therefore see why the (properly normalized) periods $\psi_{1}$ and $\psi_{2}$ defined in Eq. 7 are solutions to the second-order differential operator $L_{m_{t}^{2}}^{(2)}$ annihilating the first integral at $\epsilon=0$.
Similarly, the transformation to canonical form introduces $G_{1}$ as the solution to a third-order differential operator (related to $L_{36}$ discussed in Ref. [33]). In more detail, the DEs in Eq. 10 can be rewritten as a third-order differential operator annihilating the third integral, $I_{1,1,0,1,1,1,0,1}^{0,0,0}[(k_{1}+p_{1})^{2}]$, where the term in the square brackets is a numerator under the integral sign. The general solution for this DE is given by
@@
$\hskip-50.0pt+c_{2}\left(4m_{t}^{2}\psi_{2}^{2}+G_{1,2}\right)+c_{3}+\ldots\,$,
@@
that evaluates to a complete elliptic integral of the third kind $\Pi$ with argument either $\kappa^{2}$ or $1-\kappa^{2}$ (defined below Eq. 8), depending on the integration cycle. For example, we have that
@@
This immediately relates $G_{1}$ to either $\mathrm{\Pi}(\frac{e_{3}-e_{2}}{e_{3}-e_{1}},\kappa^{2})$ or $\mathrm{\Pi}(\frac{e_{1}-e_{2}}{e_{1}-e_{3}},1-\kappa^{2})$ depending on which period is chosen in the semi-simple part $W^{\rm ss}$ in Eq. 13.
@@
$\hskip-40.0pt+c_{2}\left(8\,G_{2,1}+4\,m_{t}^{2}\,\frac{\psi_{1}}{\Delta_{32}}\right)$
$\hskip-40.0pt+c_{3}\left(8\,G_{2,2}+4\,m_{t}^{2}\,\frac{\psi_{2}}{\Delta_{32}}\right)+\ldots\,$,
@@
where $\mathrm{Cut}_{\text{ell.}}$ denotes the cut of the elliptic sector, $f_{1}(z_{7},z_{9})$ is linear in $z_{7}$ and $z_{9}$, and $f_{2}(z_{7},z_{9})$ is quadratic in $z_{7}$ and $z_{9}$. After a change of variables, $\{z_{7},z_{9}\}\to\{x,y\}$, that rationalizes both square roots in the denominator, the integral in $x$ can be performed to show that the integrand of Eq. 24 has the form
@@
$\hskip-70.0pt+\mathrm{E}(k)\frac{(a-1)k^{\prime}-(k-1)a^{\prime}}{2\sqrt{a-1}\sqrt{a}\sqrt{a-k}(k-1)}\,$,
where the prime indicates a partial derivative w.r.t. some kinematic variable. We emphasize that the r.h.s. of Eq. 27 does not contain elliptic integrals of the third kind. For example, this relation can be used to determine the prefactor of $\Pi(a,k)$ when the arguments $a$ and $k$ have been identified. Moreover, once the result for one integration cycle is known, the replacement